We begun the day with a little work up to be sure we stay in good shape as the lack of physical
activities could prove to be a source of problem. We are not yet in SIM conditions but we try to
implement them little by little. The full SIM conditions will be reached by tomorrow morning.
An EVA took place this morning to check out a lake area that seemed promising according to our
geologist. He came back with some samples and will analyze them in the next few days.
During this time, the pictures taken last night during an observation session were processed.
The cook and another crew member prepared a detailed list of all our menu for the rest of the
rotation in order to make sure we wouldn't fall short in supplies. Some cleaning took also place at
the same time because the windy weather we encountered yesterday brought us a lot of sand back in
the Hab. The media team checked out the material collected so far and tested some new equipment.
They will begin to process the pictures and video in the next few days to provide our first video.
We headed back to Hanksville at the end of the afternoon to be briefed and came back just before
The sky is cloudy and we don't expect to be able to use the Observatory tonight.